OpenStudy (anonymous):

MEDAL and my gratitude <3 Which choice is the equation of a line that passes through the point (-2,8) and is perpendicular to the line represented by this equation 5x-2y=-10 A) 2x-5y=26 B) 5x+2y=10 C) 2x+5y=36 D)-2x+5y=-13

OpenStudy (nicholausblackmon):

@dan815

OpenStudy (princeharryyy):

do u want the answer or the solution

OpenStudy (princeharryyy):

@ardentpikachu

OpenStudy (anonymous):

answer please I'm kind in rush so I can make it to my evening class

OpenStudy (princeharryyy):

C

OpenStudy (princeharryyy):

nything else @ardentpikachu

OpenStudy (anonymous):

yeah hold up :)

OpenStudy (anonymous):

Which choice is the equation of a line that passes through the point (0,-15) and is parallel to the line represented by this equation 5x-4y=12 A) y= 4/5x-8 B) y=5/4x-15 C) y=-5/4x+7 D)y=-4/5x-12

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

b

OpenStudy (anonymous):

Which choice could be the equation of a line parallel to the line represented: y=2x-3 A) Y= -1/2x+4 B) Y=2x+7 c) Y= -2x+1 Y= 3/2x-5

OpenStudy (princeharryyy):

b

OpenStudy (princeharryyy):

y = 2x +7

OpenStudy (princeharryyy):

@ardentpikachu

OpenStudy (anonymous):

Thank you so muchh

OpenStudy (princeharryyy):

wlcm :)